If an affected female XLX marries an unaffected male XY, what is the probability of passing the hypertrichosis gene to her offspring? Remember hypertrichosis is an X-linked dominant disease.

If an affected male XLY marries an unaffected female XX, what is the probability of passing the hypertrichosis gene to his offspring?
Answer #1

If an affected female XLX marries an unaffected male XY,

Possible genetic combination of the offsprings taking one gene from male and one from female,

XLX

XLY

XX

XY

Since each type of offspring has equal probability of being born, therefore

P[ Passage of hypertrichosis gene to offspring] = 2/4 = 0.5

.

If an affected male XLY marries an unaffected female XX.

Possible genetic combination of the offsprings taking one gene from male and one from female,

XLX

XLX

XY

XY

Since each type of offspring has equal probability of being born, therefore

P[ Passage of hypertrichosis gene to offspring] = 2/4 = 0.5